# TelePhantom: A User-Friendly Teleoperation System with Virtual Assistance for Enhanced Effectiveness

### v1

@@ IV-B1 Sim-to-Real Alignment Using AprilTag
As shown in Fig. 5, our calibration procedure involves multiple sequential transformations. We begin with hand-eye calibration of the fixed top-view camera (1), followed by detecting the AprilTag’s pose from this camera’s perspective (2). Using these relationships, we can compute the transformation between the AprilTag and the robot base frame (2). For the additional floating third-view camera, we first detect the AprilTag pose from its viewpoint (3), which then allows us to establish its position relative to the robot base frame through the standard AprilTag reference (3). This chain of transformations ensures consistent phantom visualization regardless of camera movement or viewpoint changes, making our system robust for dynamic viewing scenarios.
IV-B2 Robot Phantom Visualization
Building upon this spatial alignment system, we implement the visual component of our phantom system. As shown in Fig. 6(a), we use the pyrender [31] library to render the virtual robot through the previous float camera. The rendered phantom image is then composited with the actual camera feed using alpha blending, creating a seamless overlay where the virtual robot appears naturally integrated with the physical environment. This visualization approach requires careful calibration of the virtual camera parameters to match the physical cameras, ensuring that the phantom accurately reflects the robot’s intended configuration in the workspace. The resulting overlay provides users with an intuitive preview of planned motions.
IV-B3 Multi-view Phantom Visualization
Task TelePhantom Open Teach [3] AnyTeleop[6] Telekinesis[32]
@@ IV-B3 Multi-view Phantom Visualization
TABLE I: Real Robot Teleoperation Results. For the baseline methods, we use success rate as a metric reported in their papers. Especially, we use the success rate of an expert as a metric for Open Teach [3].
(a) Real-time phantom rendering
@@
w/o phantom w/ phantom Difference $\uparrow$ w/o phantom w/ phantom Difference $\downarrow$
@@
TABLE II: Effect of Phantom Assistance on New User Performance.
Our phantom system supports multiple viewpoints to enhance spatial awareness during complex manipulation tasks. As shown in Fig. 6(b) and Fig. 6(c), we deploy cameras to capture both third-person and top-down views of the workspace, allowing users to verify planned movements from complementary perspectives. For instance, during manipulation tasks, users can simultaneously monitor the robot’s vertical position from the third-person view while checking its planar alignment from the top-down view.
IV-B4 IO Control for Phantom Visualization
Figure 7: State transition of phantom control
Our system uses a foot pedal as an intuitive control interface for state transitions. As shown in Fig. 7, when the IO is activated, the physical robot stops while the phantom appears and responds to user commands for motion preview. Upon IO deactivation, the phantom disappears, and the system extracts its final configuration as the target pose, bypassing any exploratory movements made during the preview. This target is then processed by the mplib motion planning library [33] to generate an optimal trajectory for execution. After reaching the target pose, the robot automatically resumes real-time teleoperation until the next IO is activated. This workflow ensures users execute only refined, intentional movements, eliminating exploratory motions from demonstration data.
V Experiments
@@
(Q1) How effective is TelePhantom vs Baselines?
(Q2) How beneficial is TelePhantom for new users?
(Q3) How versatile is TelePhantom on robots?
V-A Experiment Setup
TelePhantom integrates several key components. The user’s movements are captured via an inertial motion capture suit and a pair of strain gauge-based mocap gloves, providing high-fidelity joint angle data for both body posture and delicate hand movements. To visualize robot action before execution, we employ a RealSense RGB-D camera and render the phantom either on a standard 2D display or within a VR headset (e.g., Meta Quest 3). For the physical hardware, we use a UFactory xArm-6 robotic arm outfitted with a Leap hand [29], utilizing AprilTag [30] to align the initial poses of the phantom and the real robot.
@@ V-A Experiment Setup
V-B Real Robot Teleoperation
We test TelePhantom on five real-world tasks that highlight different manipulation challenges:
@@ V-B Real Robot Teleoperation
Performance Comparison with Baselines. Addressing Q1 (effectiveness vs. baselines), we compare the real robot teleoperation performance of TelePhantom, operated by an expert, with Telekinesis [32], AnyTeleop [6] and Open Teach [3]. As summarized in Table I, TelePhantom outperforms baseline methods in terms of success rate. The baseline methods are limited to simple, short-horizon tasks. While the baseline methods can handle basic tasks, they fail to accurately capture subtle wrist rotations and movements in more delicate operations like the hung task. The superior performance of TelePhantom in these scenarios demonstrates its capability to handle more challenging tasks effectively.
Effect of Phantom. To address Q2 (benefits for new users), we assess TelePhantom’s utility by measuring both the success rate and the average execution time on the real robot (demonstration time in imitation learning) with and without phantom. As shown in Table II, TelePhantom significantly improves the performance of new users by improving their task success rate. The execution time on a real robot corresponds to the demonstration time required for imitation learning. High-quality demonstrations are typically associated with shorter execution times. With the assistance of TelePhantom, new users achieve a lower average execution time, indicating that they can produce higher-quality demonstrations compared to users without access to TelePhantom. For new users, command visualization and reduced execution time not only decrease physical strain but also improve safety.
V-C Deployment on Various Robots
TelePhantom’s modular design enables seamless adaptation to different input devices and end-effectors. Thanks to our robot-centric phantom approach and standardized transformation framework, users can readily configure the system for various robotic platforms by simply updating the kinematic parameters and transformation chain. This adaptability across different robot configurations, as posed in Q3, demonstrates TelePhantom’s versatility as a flexible teleoperation solution that can serve diverse robotic applications with minimal modification requirements.
VI Conclusion
@@
In this paper, we presented TelePhantom, a low-cost teleoperation system featuring AR phantom feedback designed for deployment across diverse robotic platforms. Our experiments demonstrate that TelePhantom offers sufficient flexibility and safety to successfully perform a variety of fine-grained tasks. Through its innovative phantom visualization approach, the system provides immediate, robot-centric feedback that enables users to preview and refine their commands before physical execution.
@@
While our results show promising advances in teleoperation interfaces, a key limitation is the visual ambiguity caused by occlusions between the phantom and scene objects. Future work could address this by incorporating depth information from RGB-D cameras to create more accurate spatial relationships between virtual and physical elements.
Addressing these visualization challenges could yield several essential benefits. Improved occlusion handling would reduce user uncertainty during complex manipulation tasks, leading to higher-quality demonstration data for imitation learning. Additionally, more accurate spatial rendering would create a more intuitive teleoperation experience, potentially reducing user training time and cognitive load. These improvements would further advance TelePhantom’s goal of providing accessible, effective teleoperation capabilities across diverse robotic applications.

### v4

@@ V-B1 Spatial Alignment Using AprilTag
As shown in Fig. 5, our calibration procedure involves multiple sequential transformations. We begin with hand-eye calibration of the fixed top-view camera (1), followed by detecting the AprilTag’s pose from this camera’s perspective (2). Using these relationships, we can compute the transformation between the AprilTag and the robot base frame (2). For the additional floating third-view camera, we first detect the AprilTag pose from its viewpoint (3), which then allows us to establish its position relative to the robot base frame through the standard AprilTag reference (3). This chain of transformations ensures consistent preview visualization regardless of camera movement or viewpoint changes, making our system robust for dynamic viewing scenarios.
V-B2 Visual Preview Visualization
Building upon this spatial alignment system, we implement the visual component of our preview system. As shown in Fig. 6(a), we use the pyrender [35] library to render the virtual robot through the previous float camera. The rendered preview image is then composited with the actual camera feed using alpha blending, creating a seamless overlay where the virtual robot appears naturally integrated with the physical environment. This visualization approach requires careful calibration of the virtual camera parameters to match the physical cameras, ensuring that the preview accurately reflects the robot’s intended configuration in the workspace. The resulting overlay provides users with an intuitive preview of planned motions. Our preview system also supports multiple viewpoints to enhance spatial awareness during complex manipulation tasks. As shown in Fig. 6(b) and Fig. 6(c), we deploy cameras to capture both third-person and top-down views of the workspace, allowing users to verify planned movements from complementary perspectives.
Task TelePreview Open Teach [3] AnyTeleop[6] Telekinesis[36]
@@ V-B2 Visual Preview Visualization
TABLE I: Real Robot Teleoperation Results. We use the success rates the baseline methods report in their papers, marking tasks they did not attempt or do not support with “-”. Success rates for Open Teach [3] reflect expert performance.
(a) Preview Rendering
@@
w/o preview w/ preview Difference $\uparrow$ w/o preview w/ preview Difference $\downarrow$
@@
TABLE II: Effect of Preview Assistance on New User Performance.
V-B3 IO Control for Preview Visualization
Our system uses a foot pedal as an intuitive control interface for state transitions. As shown in Fig. 7, when the IO is activated, the physical robot stops while the preview appears and responds to user commands for motion preview (Preview Mode). Upon IO deactivation, the preview disappears, and the system extracts its final configuration as the target pose to align, bypassing any exploratory movements made during the preview (Align Mode). This target is then processed by the mplib motion planning library [37] to generate an optimal trajectory for execution. After reaching the target pose, the robot automatically resumes real-time teleoperation until the next IO is activated. This workflow ensures users execute only refined, intentional movements, eliminating exploratory motions from demonstration data.
Figure 7: State Transition of Preview Control
VI Experiments
@@
Q1: How effective is TelePreview vs. Baselines?
Q2: How beneficial is TelePreview for new users?
Q3: How adaptable is TelePreview for new hardware?
VI-A Experiment Setup
TelePreview integrates several key components. We capture the user’s movements via an inertial motion capture suit [38] and a pair of strain-gauge-based mocap gloves [33], providing high-fidelity joint angle data for both body posture and delicate hand movements. To visualize robot action before execution, we employ a RealSense RGB-D camera and render the preview either on a standard 2D display or within a VR headset (e.g., Meta Quest 3[39]). For the physical hardware, we use a UFactory xArm-6 robotic arm[40] outfitted with a Leap hand [34], utilizing AprilTag [31] to align the initial poses of the preview and the real robot.
@@ VI-A Experiment Setup
VI-B Task Descriptions
We test TelePreview on five real-world tasks, each highlighting different manipulation challenges:
@@ VI-B Task Descriptions
VI-C Evaluation Metrics
We measure two key metrics for each task:
  • Success Rate: We define the success rate as the proportion of times the operator completes the task within the given constraints. For example, if the user attempts a cup-hanging task $X$ times and successfully hangs the cup on the rack $Y$ times, the success rate is $\frac{Y}{X}$.
  • Execution Time: The total duration from task initiation to successful completion (i.e., the length of the recorded demo episode).
VI-D Performance Comparison with Baselines (Q1)
To address Q1—evaluating how our approach compares to existing teleoperation methods—we measure the actual robot performance of TelePreview against three baselines: Telekinesis [36], AnyTeleop [6], and Open Teach [3]. For fairness and consistency, an expert user operates each method and tests it on the five tasks described in Section VI-B. The primary metrics are success rate and average execution time.
As summarized in Table I, TelePreview achieves higher success rates than all baselines. While Telekinesis, AnyTeleop, and Open Teach handle simpler, short-horizon tasks, they often struggle with complex manipulations requiring subtle wrist rotations and fine-grained positioning (e.g., the Hang task). TelePreview, by contrast, captures these nuances more effectively, reflecting stronger overall performance and versatility. These findings demonstrate TelePreview’s ability to address a broad spectrum of real-world challenges, from basic pick-and-place operations to more intricate tasks involving delicate movements.
(a) LEAP Hand
(b) Gripper
Figure 8: Deployment on Different Robots.
VI-E Effect of Preview for Users (Q2)
To investigate Q2—how the preview feature benefits new (non-expert) users—we evaluate both success rate and average execution time in two conditions: (1) with the preview enabled and (2) with the preview disabled. The same five tasks from Section VI-B are attempted by new users, with each task repeated 10 times.
As shown in Table II, enabling the preview boosts user performance. Success rates increase across all tasks, with especially notable gains in challenging manipulations such as cup stacking (+0.5) and pick & place (+0.4). Moreover, average execution times drop by up to 12.63 seconds (cup stacking), demonstrating not only improved accuracy but also greater efficiency. By visualizing a virtual “preview” of the robot’s intended motion and confirming the final pose prior to executing physical movements, users eliminate the need for exploratory motions in the actual workspace. This design ensures that recorded trajectories contain only intentional, task-oriented actions without extraneous trial-and-error, providing cleaner, high-quality demonstrations for imitation learning. As a result, TelePreview delivers both a more intuitive teleoperation experience and more consistent data for downstream training.
VI-F Deployment on Various Robots (Q3)
We deploy our TelePreview system on a Ufactory xArm [40] equipped with both a LEAP hand [34] and a parallel gripper. By updating only the kinematic parameters and transformation chain for each end-effector, TelePreview preserves its full functionality across diverse setups, as shown in Fig. 8. This successful deployment on multiple end-effector configurations illustrates the system’s adaptability to new hardware. Its modular design allows users to seamlessly integrate TelePreview into different robotic platforms while retaining all key features—preview, multi-view visualization, and intuitive control interfaces.
VII Conclusion
@@
In this paper, we presented TelePreview, a low-cost teleoperation system featuring AR preview feedback designed for deployment across diverse robotic platforms. Our experiments demonstrate that TelePreview offers sufficient flexibility and safety to successfully perform a variety of fine-grained tasks. TelePreview provides immediate, robot-centric feedback that enables users to preview and refine their commands before physical execution.
@@
While our results show promising advances in teleoperation interfaces, a key limitation is the visual ambiguity caused by occlusions between the preview robot and scene objects. Future work could address this by incorporating depth information from RGB-D cameras to create more accurate spatial relationships between virtual and physical elements.